Good Morning
Help Word problems so lost
A store mixes Kentucky bluegrass worth $11 per pound and chewing fesue worth $15 per pound mixture to sell for $12 per pound. How much of each is to be used to make a 276Ld.
Thank you betty

Answer by Alan3354(69443)

Same problem, different numbers.
===================================================
If you have a solution that is .25 orange juice and a solution that is 100% pure orange juice how much of each to get 17 liters that is 60% orange juice.
====
.25 = 25%
Do it like this one.
----
One canned Juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together to get 10L that is 6% orange juice.
---
t = 15%
f = 5%
---
t + f = 10 (total solution)
15t + 5f = 10*6 (total OJ)
Solve for t & f